When Mr Smith’s monthly income was €5,000, he usually went to the local coffee shop 12 times a month. Mr Smith had a pay rise and his current income is €5,500 a month. Now, he goes to the local coffee shop 10 times a month. Compute the income elasticity of demand using the midpoint method. Explain your answer. Is having a coffee at the local coffee shop normal or inferior good to Mr Smith? Illustrate your answer using an appropriate graph.
1
Expert's answer
2019-12-05T09:39:17-0500

The income elasticity of demand using the midpoint method is:

Ei = (10 - 12)/(5,500 - 5,000)×(5,500 + 5,000)/(10 + 12) = -2/500×10,500/22 = -1.91, so a coffee at the local coffee shop is inferior good to Mr Smith.